An a.c. voltage of 5Vr.m.s and frequency 200 Hz is connected in series to a 1100 Ω resistor, 0.012 H inductor and 8μF capacitor. Calculate the reactance of the inductor.
1
Expert's answer
2020-07-20T14:55:51-0400

The reactance of the inductor is givem by the following expression:


"X_L = 2\pi fL"

where "f = 200 Hz" the frequency and "L = 0.012H" is the inductance. Tus, obtain the following numbers:


"X_L = 2\pi\cdot 200\cdot 0.012 \approx 150.8 \Omega"

Answer. 150.8 Ohm.